OpenStudy (anonymous):

Consider the combustion of carbon monoxide (CO) in oxygen gas: 2CO(g) + O2 (g) --> 2CO2(g) Starting with 3.84 moles of CO, calculate the number of moles of CO2 produced if there is enough oxygen gas to react with all the CO.

OpenStudy (aaronq):

Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) and solve for the moles of \(CO_2\), (\(n_{CO_2}\)).